OpenStudy (anonymous):

why does (n+1)n ^{n} / (n+1)^{n+1} simplify down to (n ^{n})/ (n+1^{n})

OpenStudy (anonymous):

exponent laws, simply :  what I did do with the denominator? Here it is : (n+1)^(n+1) = (n+1)^1 * (n+1)^n (rule being a^n*a^m = a^(m+n))